General Physics

what are the types of combination of capacitor/

puja kumari
12 Years agoGrade

1 Answer

Apoorva Arora
11 Years ago
Some circuits have capacitors connected in series and parallel combinations. To determine, for instance, the total charge stored by the set of capacitors, we have to find the single equivalent capacitance of the set. This is done by identifying a pair of capacitors in the set that are either in series or in parallel with one another, replacing that pair with its equivalent capacitor (thereby reducing the number of capacitors by one) and iterating until we're left with one capacitor that is the equivalent of the set.
for parallel combination
C_{eq}=C_{1}+C_{2}+C_{3}.....
For series combination
\frac{1}{C_{eq}}=\frac{1}{C_{1}}+\frac{1}{C_{2}}+\frac{1}{C_{3}}....